# Large Language Model Inference Acceleration: A Comprehensive Hardware Perspective

### v3

@@ 3.2.1. Overview
Figure 5. Sparsity and sparsity patterns.
Table 4. Sparsity on CPU, GPU, FPGA, ASIC, and PIM/NDP
@@ 3.2.2. CPU
Activation Sparsity. Activation sparsity is determined by activation functions. Commonly using SwiGLU (shazeer2020glu, ) and GeGLU (hendrycks2016gaussian, ) exhibits limited sparsity for LLMs, but simply replacing these functions with ReLU fails to achieve sufficient sparsity. Turbo Sparse (song2024turbo, ) proposes the dReLU activation function to improve LLM activation sparsity, along with a high-quality training data mixture ratio to facilitate effective sparsity. By applying their sparsity method to the Mistral and Mixtral models, only 2.5 billion (35.7%) and 4.3 billion (9.2%) parameters are activated per inference iteration, respectively. For Mistral-7B, Turbo Sparse achieves 8.71 tokens/s and 9.94 tokens/s on Intel i9-14900HX processor and Intel i7-12700K processor, respectively. For Mixtral-47B with 4-bit quantization, Turbo Sparse achieves 16.1 tokens/s, 11.98 tokens/s and 11.1 tokens/s on Intel i9-14900HX, Intel i7-12700K and SnapDragon 8 Gen3, respectively. ProSparse (song2024prosparse, ) also introduces activation function substitution, progressive sparsity regularization, and activation threshold shifting to help non-ReLU LLMs obtain high activation sparsity without performance degradation. For Llama2-7B and Llama2-13B, ProSparse achieves high sparsity of 89.32% and 88.80%, and 16.3 tokens/s and 8.67 tokens/s, respectively, based on PowerInfer (song2023powerinfer, ) framework.
3.2.3. GPU
@@
Weight Sparsity. LLM-pruner (ma2023llm, ), adopts structural pruning that selectively removes non-critical coupled structures based on gradient information, maximally preserving the majority of the LLM’s functionality. To this end, the performance of pruned models can be efficiently recovered through tuning techniques, LoRA, in merely 3 hours, requiring only 50K data. We validate the LLM-Pruner on three LLMs, including Llama, Vicuna, and ChatGLM, and demonstrate that the compressed models still exhibit satisfactory capabilities in zero-shot classification and generation.
@@
LLMs can be pruned to at least 50% sparsity in one-shot, without any retraining, at minimal loss of accuracy. SparseGPT (frantar2023sparsegpt, ) requires a sophisticated weight update procedure in an iterative pruning process. Wanda (sun2023wanda, ) prunes weights with the smallest magnitudes multiplied by the corresponding input activations, on a per-output basis. Notably, Wanda requires no retraining or weight update, where pruning process is faster. Besides unstructured pattern, these two methods generalizes to semi-structured N:M (2:4 and 4:8) patterns. E-Sparse (li2023sparse, ) introduces entropy to quantify the information richness within each channel (intra-channel) of the input features, and adopts it to enhance the feature norms (crosschannel) as a metric to evaluate parameter importance. Furthermore, it proposes Channel Shuffle to reorder the information distribution in LLMs to obtain N:M Sparsity with less information loss. 2:4 sparsity as supported by NVIDIA GPUs of generation Ampere and newer theoretically offers 2$\times$ acceleration of matrix multiplications. In practical, 2:4 sparsity can achieve 1.54$\times$-1.79$\times$ speedup for MatMul, and end-to-end speedups are about 1.21$\times$-1.25$\times$ (due to some extra overheads from e.g. attention).
@@
Based on the key observation that the bottleneck of LLM inference is the skinny matrix multiplications, Flash-LLM (xia2023flash, ) proposes a general Load-as-Sparse and Compute-as-Dense methodology for unstructured sparse matrix multiplication. Flash-LLM proposes a new sparse format called Tiled-CSL to relieve the memory bandwidth bottleneck and support the tile-by-tile SpMM execution with tensor cores. For OPT-30B, Flash-LLM achieves 80% sparsity with 1.44% accuracy decrease and about 290 tokens/s, 500 tokens/s, 800 tokens/s, and 1187 tokens/s on single A100 GPU with batch sizes 8, 16, 32, and 64, respectively.
@@
Agarwalla et al. (agarwalla2024enabling, ) combine the SparseGPT one-shot pruning method and sparse pretraining to pretrain a high sparsity LLM. They deploy model on GPU and CPU by utilizing Neural Magic’s DeepSparse engine and Neural Magic’s nm-vllm engine, respectively. For Llama-7B, on NVIDIA A10 GPU, they achieve 44.4 tokens/s and 47.9 tokens/s with 50% sparsity and 70% sparsity, respectively. On AMD EPYC 9R14 Processor, they achieve 4.4 tokens/s and 6.9 tokens/s with 50% sparsity and 70% sparsity, respectively.
@@
Existing methods require costly retraining, forgo LLM’s in-context learning ability, or do not yield wall-clock time speedup on modern hardware. DejaVu (liu2023deja, ) predicts contextual sparsity on the fly given inputs to each layer, along with an asynchronous and hardware-aware implementation that speeds up LLM inference. For OPT-175B model, DejaVu achieves up to 75% sparsity and 50 tokens/s on 8 A100-80GB GPUs with batch size 1, which is over 2$\times$ and 6$\times$ faster than FasterTransformer and Hugging Face implementation, respectively.
@@
Attention Sparsity. During the prefilling phase of LLM inference, attention computation complexity scales quadratically with input sequence length. Given limited GPU computing and memory resources, attention sparsification can reduce the number of attention values to accelerate prefilling phase. For static sparsity, Sparse Transformer (child2019generating, ), StreamingLLM (xiao2023streamingllm, ), Bigbird (zaheer2020big, ), and Longformer (beltagy2020longformer, ) use the manual combination of global and local patterns to replace the full attention patterns. The local pattern captures the local context of each token within a fixed size or stride while the global pattern captures the relationship between the specific tokens to all other tokens. For Llama2-7B and Llama2-13B models, StreamingLLM achieves 15.38-32.26 tokens/s and 9.43-20.83 tokens/s on single NVIDIA A6000 GPU, respectively. For dynaimic sparsity, Adaptively Sparse Attention (correia2019adaptively, ) replaces softmax with $\alpha$-entmax, a differentiable generalization of softmax that allows low-scoring words to receive precisely zero weight and drops parts of the context that are no longer required for future generation. Reformer (kitaev2020reformer, ) replaces dot-product attention by using locality-sensitive hashing, changing the complexity from $O(L^{2})$ to $O(LlogL)$, where $L$ is the sequence length. Sparse Flash Attention (pagliardini2023faster, ) extends FlashAttention (dao2022flashattention, ) GPU kernel and encompasses key/query dropping and hashing-based attention. Sparse Sinkhorn Attention (tay2020sparse, ) adopts a learned sorting network to align keys with their relevant query buckets, ensuring that attention is computed only between the corresponding query-key pairs. H2(zhang2024h2o, ) observes that a small portion of tokens (called Heavy Hitters, H2) contributes most of the value when computing attention scores. H2O introduces a dynamic attention sparsification method to adopt KV cache eviction policy that dynamically retains a balance of recent and H2 tokens. For OPT-6.7B model, H2O with 20% H2 achieves 30.4 tokens/s on single NVIDIA T4 GPU.
3.2.4. FPGA
@@
Weight Sparsity. FlightLLM (zeng2024flightllm, ) is the first real FPGA-based LLM accelerator which proposes a configurable sparse DSP chain to support different sparsity patterns with high computation efficiency. Then, it proposes an always-on-chip decode scheme to boost memory bandwidth with mixed-precision support. Finally, it proposes a length adaptive compilation method to reduce the compilation overhead. For Llama2-7B model, FlightLLM achieves 55 tokens/s and 92.5 tokens/s with batch size 1 on the Xilinx Alveo U280 FPGA and Versal VHK158 FPGA, respectively. EdgeLLM (huang2024edgellm, ) integrates 4-bit weight-only quantization and utilizes log-scale structural sparsity for weight parameters in the matrix multiplication operator. For ChatGLM2-6B model, it achieves average 67 tokens/s with 55.07W power consumption on AMD Xilinx VCU128 FPGA.
3.2.5. ASIC
@@
Attention Sparsity. Spatten (wang2021spatten, ) leverages token sparsity, head sparsity, and quantization opportunities to reduce the attention computation and memory access. It assesses the cumulative importance of each word by aggregating the attention matrix columns, subsequently pruning tokens with minimal cumulative significance from the input in subsequent layers. It provides 2.88 TOPS computing power and 8.3W power consumption by considering all memory access under 40nm. Experimental results shows that for GPT2-Medium, it can achieve about 35.86 tokens/s in decode stage. TF-MVP (yoo2023tf, ) quantitatively analyzes sparsity patterns of pruned-transformer models with the cutting-edge fine-grained pruning scheme for the first time and presents the mixed-length vector pruning (MVP) procedure by utilizing this direction strength. From hardware perspective, it introduces the TF-MVP architecture, a sparsity-aware cost-efficient accelerator design dedicated to the proposed pruned-transformer models. Implemented in a 28nm CMOS technology at 400MHz, TF-MVP provides 0.835 TOPS with 1.721W on-chip power consumption for accelerating GPT-2 small model. SOFA (wang2024sofa, ) predicts attention sparsity by using log-based add-only operations to avoid the significant overhead of prediction. Then, a distributed sorting and a sorted updating FlashAttention mechanism are proposed with a cross-stage coordinated tiling principle, which enables fine-grained and lightweight coordination among stages, helping optimize memory access and latency. SOFA provides 24.423 TOPS computing power and 3.4W power consumption under 28nm. For Llama2-7B, compared to NVIDIA A100 GPU, it achieves 9.5$\times$ inference speedup in prefill stage.
3.2.6. PIM/NDP
@@
Weight Sparsity. LauWS (li2024lauws, ) proposes an unstructured sparsity method for NDP systems and is evaluated on a practical GDDR6-based bank NDP. Not only the overall features of the total matrix but the features of the local region are preserved by ignoring non-feature values as much as possible, which is beneficial to the trade-off between high sparsity and least accuracy loss. Compared to dense models, it achieves 1.23$\times$ and 1.24$\times$ speedup for GPT-2 small model and OPT-125M at 50% sparsity, respectively. Sharda et al. (sharda2024accelerator, ) propose to use the capacitorless 3D stackable DRAM to store much larger LLMs compared to conventional DRAM at higher density. To reduce the intermediate data size, they propose to use a layer-wise sparsity-quantization hybrid (LSQH) algorithm, which induces sparsity based on calculations performed using low-bit quantization to reduce both the energy consumption and the data storage requirements. Finally, a 3D heterogeneously integrated accelerator is designed by stacking a 3D DRAM with logic dies designed in the 3nm technology node at 1GHz. The evaluation shows that for Llama2-13B, it achieves 163k tokens/s in prefill stage with 193W power consumption.
@@
Attention Sparsity. HARDSEA (liu2023hardsea, ) proposes an attention sparsity method by predicting lightweight token relevance and design a hybrid analog-ReRAM and digital-SRAM in-memory computing accelerator. It employs ReRAM-CIM, whose precision is sensitive to circuit non-idealities, to take charge of token relevance prediction where only computing monotonicity is demanded. The SRAM-CIM, utilized for exact sparse attention computing, is reorganized as an on-memory-boundary computing scheme, thus adapting to irregular sparsity patterns. Experimental results show that HARDSEA prunes BERT and GPT-2 small model to 20% sparsity without accuracy loss, achieving 5.8$\times$–6.7$\times$ speedup over NVIDIA RTX 3090 GPU.
@@
Figure 6. LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 1) vs power on different platforms with sparsity.
3.2.7. Comparison

### v4

@@ 3.2.1 Overview
Figure 5: Sparsity and sparsity patterns.
Table 4: Sparsity on CPU, GPU, FPGA, ASIC, and PIM/NDP
@@ 3.2.2 CPU
Activation Sparsity. Activation sparsity is determined by activation functions. Commonly using SwiGLU [290] and GeGLU [291] exhibits limited sparsity for LLMs, but simply replacing these functions with ReLU fails to achieve sufficient sparsity. Turbo Sparse [171] proposes the dReLU activation function to improve LLM activation sparsity, along with a high-quality training data mixture ratio to facilitate effective sparsity. By applying their sparsity method to the Mistral and Mixtral models, only 2.5 billion (35.7%) and 4.3 billion (9.2%) parameters are activated per inference iteration, respectively. For Mistral-7B, Turbo Sparse achieves 8.71 tokens/s and 9.94 tokens/s on Intel i9-14900HX processor and Intel i7-12700K processor, respectively. For Mixtral-47B with 4-bit quantization, Turbo Sparse achieves 16.1 tokens/s, 11.98 tokens/s and 11.1 tokens/s on Intel i9-14900HX, Intel i7-12700K and SnapDragon 8 Gen3, respectively. ProSparse [172] also introduces activation function substitution, progressive sparsity regularization, and activation threshold shifting to help non-ReLU LLMs obtain high activation sparsity without performance degradation. For Llama2-7B and Llama2-13B, ProSparse achieves high sparsity of 89.32% and 88.80%, and 16.3 tokens/s and 8.67 tokens/s, respectively, based on PowerInfer [259] framework.
Weight Sparsity. SparAMX [173] utilizes Advanced Matrix Extensions (AMX) support on the latest Intel CPUs with unstructured sparsity in linear layers to achieve a 1.42× reduction in end-to-end latency compared to the PyTorch implementation. It also provides a set of open-source customized sparse kernels to support all linear layers.
3.2.3 GPU
Activation Sparsity. SoLA [192] leverages Soft activation sparsity and Low-rank decomposition. SoLA can identify and retain a minority of components significantly contributing to inference, while compressing the majority through low-rank decomposition, based on the activation pattern in the FFNs of LLMs. For Llama2-7B and 13B models, it achieves 76.69 tokens/s and 55.96 tokens/s on RTX 4090 GPU, respectively. R-Sparse [193] finds two key observations of FFNs: (i) the nonsparse components of the input can be regarded as a few bias terms, and (ii) the full computation can be effectively approximated by an appropriate combination of input channels and weight singular values. So it replaces the linear layers in LLMs with a rank-aware sparse inference method that leverages the sparsity of input channels and singular value components. For the Llama2-7B and Llama3-8B models, it achieves around 20 tokens/s and 18 tokens/s on a NVIDIA A6000 GPU, respectively.
@@ 3.2.3 GPU
Weight Sparsity. LLM-pruner [174], adopts structural pruning that selectively removes non-critical coupled structures based on gradient information, maximally preserving the majority of the LLM’s functionality. To this end, the performance of pruned models can be efficiently recovered through tuning techniques, LoRA, in merely 3 hours, requiring only 50K data. We validate the LLM-Pruner on three LLMs, including Llama, Vicuna, and ChatGLM, and demonstrate that the compressed models still exhibit satisfactory capabilities in zero-shot classification and generation.
@@
LLMs can be pruned to at least 50% sparsity in one-shot, without any retraining, at minimal loss of accuracy. SparseGPT [175] requires a sophisticated weight update procedure in an iterative pruning process. Wanda [176] prunes weights with the smallest magnitudes multiplied by the corresponding input activations, on a per-output basis. Notably, Wanda requires no retraining or weight update, where pruning process is faster. Besides unstructured pattern, these two methods generalizes to semi-structured N:M (2:4 and 4:8) patterns. E-Sparse [177] introduces entropy to quantify the information richness within each channel (intra-channel) of the input features, and adopts it to enhance the feature norms (crosschannel) as a metric to evaluate parameter importance. Furthermore, it proposes Channel Shuffle to reorder the information distribution in LLMs to obtain N:M Sparsity with less information loss. 2:4 sparsity as supported by NVIDIA GPUs of generation Ampere and newer theoretically offers 2$\times$ acceleration of matrix multiplications. In practical, 2:4 sparsity can achieve 1.54$\times$-1.79$\times$ speedup for MatMul, and end-to-end speedups are about 1.21$\times$-1.25$\times$ (due to some extra overheads from e.g. attention).
@@
Based on the key observation that the bottleneck of LLM inference is the skinny matrix multiplications, Flash-LLM [178] proposes a general Load-as-Sparse and Compute-as-Dense methodology for unstructured sparse matrix multiplication. Flash-LLM proposes a new sparse format called Tiled-CSL to relieve the memory bandwidth bottleneck and support the tile-by-tile SpMM execution with tensor cores. For OPT-30B, Flash-LLM achieves 80% sparsity with 1.44% accuracy decrease and about 290 tokens/s, 500 tokens/s, 800 tokens/s, and 1187 tokens/s on single A100 GPU with batch sizes 8, 16, 32, and 64, respectively.
@@
Agarwalla et al. [179] combine the SparseGPT one-shot pruning method and sparse pretraining to pretrain a high sparsity LLM. They deploy model on GPU and CPU by utilizing Neural Magic’s DeepSparse engine and Neural Magic’s nm-vllm engine, respectively. For Llama-7B, on NVIDIA A10 GPU, they achieve 44.4 tokens/s and 47.9 tokens/s with 50% sparsity and 70% sparsity, respectively. On AMD EPYC 9R14 Processor, they achieve 4.4 tokens/s and 6.9 tokens/s with 50% sparsity and 70% sparsity, respectively.
@@
Existing methods require costly retraining, forgo LLM’s in-context learning ability, or do not yield wall-clock time speedup on modern hardware. DejaVu [180] predicts contextual sparsity on the fly given inputs to each layer, along with an asynchronous and hardware-aware implementation that speeds up LLM inference. For OPT-175B model, DejaVu achieves up to 75% sparsity and 50 tokens/s on 8 A100-80GB GPUs with batch size 1, which is over 2$\times$ and 6$\times$ faster than FasterTransformer and Hugging Face implementation, respectively.
For Mamba models, shihab et al. [190] propose a novel unstructured pruning framework tailored for Mamba models that achieves up to 70% parameter reduction while retaining over 95% of the original performance. They integrate three key innovations: a gradient-aware magnitude pruning technique, an iterative pruning schedule, and a global pruning strategy. As a result, they achieve 2083.33 tokens/s and 1219.51 tokens/s on both Mamba-130M and Mamba-370M models on a single NVIDIA A100 GPU.
@@
Attention Sparsity. During the prefilling phase of LLM inference, attention computation complexity scales quadratically with input sequence length. Given limited GPU computing and memory resources, attention sparsification can reduce the number of attention values to accelerate prefilling phase. For static sparsity, Sparse Transformer [181], StreamingLLM [183], Bigbird [182], and Longformer [184] use the manual combination of global and local patterns to replace the full attention patterns. The local pattern captures the local context of each token within a fixed size or stride while the global pattern captures the relationship between the specific tokens to all other tokens. For Llama2-7B and Llama2-13B models, StreamingLLM achieves 15.38-32.26 tokens/s and 9.43-20.83 tokens/s on single NVIDIA A6000 GPU, respectively. For dynaimic sparsity, Adaptively Sparse Attention [185] replaces softmax with $\alpha$-entmax, a differentiable generalization of softmax that allows low-scoring words to receive precisely zero weight and drops parts of the context that are no longer required for future generation. Reformer [186] replaces dot-product attention by using locality-sensitive hashing, changing the complexity from $O(L^{2})$ to $O(LlogL)$, where $L$ is the sequence length. Sparse Flash Attention [187] extends FlashAttention [292] GPU kernel and encompasses key/query dropping and hashing-based attention. Sparse Sinkhorn Attention [188] adopts a learned sorting network to align keys with their relevant query buckets, ensuring that attention is computed only between the corresponding query-key pairs. H2[189] observes that a small portion of tokens (called Heavy Hitters, H2) contributes most of the value when computing attention scores. H2O introduces a dynamic attention sparsification method to adopt KV cache eviction policy that dynamically retains a balance of recent and H2 tokens. For OPT-6.7B model, H2O with 20% H2 achieves 30.4 tokens/s on single NVIDIA T4 GPU.
3.2.4 FPGA
@@
Weight Sparsity. FlightLLM [133] is the first real FPGA-based LLM accelerator which proposes a configurable sparse DSP chain to support different sparsity patterns with high computation efficiency. Then, it proposes an always-on-chip decode scheme to boost memory bandwidth with mixed-precision support. Finally, it proposes a length adaptive compilation method to reduce the compilation overhead. For Llama2-7B model, FlightLLM achieves 55 tokens/s and 92.5 tokens/s with batch size 1 on the Xilinx Alveo U280 FPGA and Versal VHK158 FPGA, respectively. EdgeLLM [134] integrates 4-bit weight-only quantization and utilizes log-scale structural sparsity for weight parameters in the matrix multiplication operator. For ChatGLM2-6B model, it achieves average 67 tokens/s with 55.07W power consumption on AMD Xilinx VCU128 FPGA.
Recent work mainly assumes structured model pruning and does not work well for unstructured sparsity. The reason behind is that it is difficult to exploit data reuse from the unstructured sparsity, leaving hardware underutilized. ChatOPU [194] proposes an FPGA-based overlay processor for LLMs, to support unstructured model pruning with better data reuse. It proposes a new diagonal dataflow on a systolic array to obtain efficient data reuse for both sparse and dense matrix multiplication. Then, it develops efficient encoding and decoding for the sparse parameters to save off-chip memory traffic and boosts the off-chip bandwidth utilization with pinned on-chip KVcache allocation and coalesced access throughout the LLM inference. Experimental results show that ChatOPU on Xilinx U200 FPGA can achieve 43.21 tokens/s and 166.26 tokens/s with 1 and 4 ChatOPU core for OPT-350M. Zhang et al. [195] propose a sparse FPGA accelerator based on Fine-grained Structured Sparsity (FSS) and Channel Redistributed Fine-grained Structured Sparsity (CRFSS). For OPT-6.7B model, their approach achieves 51.744 tokens/s and 69.776 tokens/s with low-bit quantization on Alveo U280. AccLLM [145] combines 2:4 semi-structured pruning, W2A8KV4 quantization, and sparse attention to enhance computational efficiency. Extensive experimental results demonstrate for Llama2-7B model, it achieves 164 tokens/s with only 33W on Alveo U280.
3.2.5 ASIC
@@
Attention Sparsity. Spatten [196] leverages token sparsity, head sparsity, and quantization opportunities to reduce the attention computation and memory access. It assesses the cumulative importance of each word by aggregating the attention matrix columns, subsequently pruning tokens with minimal cumulative significance from the input in subsequent layers. It provides 2.88 TOPS computing power and 8.3W power consumption by considering all memory access under 40nm. Experimental results shows that for GPT2-Medium, it can achieve about 35.86 tokens/s in decode stage. TF-MVP [197] quantitatively analyzes sparsity patterns of pruned-transformer models with the cutting-edge fine-grained pruning scheme for the first time and presents the mixed-length vector pruning (MVP) procedure by utilizing this direction strength. From hardware perspective, it introduces the TF-MVP architecture, a sparsity-aware cost-efficient accelerator design dedicated to the proposed pruned-transformer models. Implemented in a 28nm CMOS technology at 400MHz, TF-MVP provides 0.835 TOPS with 1.721W on-chip power consumption for accelerating GPT-2 small model. SOFA [198] predicts attention sparsity by using log-based add-only operations to avoid the significant overhead of prediction. Then, a distributed sorting and a sorted updating FlashAttention mechanism are proposed with a cross-stage coordinated tiling principle, which enables fine-grained and lightweight coordination among stages, helping optimize memory access and latency. SOFA provides 24.423 TOPS computing power and 3.4W power consumption under 28nm. For Llama2-7B, compared to NVIDIA A100 GPU, it achieves 9.5$\times$ inference speedup in prefill stage.
3.2.6 PIM/NDP
@@
Weight Sparsity. LauWS [199] proposes an unstructured sparsity method for NDP systems and is evaluated on a practical GDDR6-based bank NDP. Not only the overall features of the total matrix but the features of the local region are preserved by ignoring non-feature values as much as possible, which is beneficial to the trade-off between high sparsity and least accuracy loss. Compared to dense models, it achieves 1.23$\times$ and 1.24$\times$ speedup for GPT-2 small model and OPT-125M at 50% sparsity, respectively. Sharda et al. [165] propose to use the capacitorless 3D stackable DRAM to store much larger LLMs compared to conventional DRAM at higher density. To reduce the intermediate data size, they propose to use a layer-wise sparsity-quantization hybrid (LSQH) algorithm, which induces sparsity based on calculations performed using low-bit quantization to reduce both the energy consumption and the data storage requirements. Finally, a 3D heterogeneously integrated accelerator is designed by stacking a 3D DRAM with logic dies designed in the 3nm technology node at 1GHz. The evaluation shows that for Llama2-13B, it achieves 163k tokens/s in prefill stage with 193W power consumption.
FlexCiM [201] presents a flexible layer-wise N:M sparsity accounting for the presence and distribution of outliers. FlexCiM enables support for diverse sparsity patterns by partitioning a digital CiM (DCiM) macro into smaller sub-macros, delivering up to 1.75× lower inference latency and 1.5× lower energy consumption compared to existing sparse accelerators.
@@
Attention Sparsity. HARDSEA [200] proposes an attention sparsity method by predicting lightweight token relevance and design a hybrid analog-ReRAM and digital-SRAM in-memory computing accelerator. It employs ReRAM-CIM, whose precision is sensitive to circuit non-idealities, to take charge of token relevance prediction where only computing monotonicity is demanded. The SRAM-CIM, utilized for exact sparse attention computing, is reorganized as an on-memory-boundary computing scheme, thus adapting to irregular sparsity patterns. Experimental results show that HARDSEA prunes BERT and GPT-2 small model to 20% sparsity without accuracy loss, achieving 5.8$\times$–6.7$\times$ speedup over NVIDIA RTX 3090 GPU.
@@
Figure 6: LLM ($\sim$ 7 billion parameters) decode stage throughput (batch size 1) vs power on different platforms with sparsity.
3.2.7 Comparison